# Creative Writers’ Attitudes on Writing as Training Data for Large Language Models

### v1

@@ Abstract.
The use of creative writing as training data for large language models (LLMS) is highly contentious. While some argue that such use constitutes “fair use” and therefore does not require consent or compensation, others argue that consent and compensation is the morally correct approach. In this paper, we seek to understand how creative writers reason about the real or hypothetical use of their writing as training data and under what conditions, if any, they would consent to their writing being used. We interviewed 33 writers with variation across genre, method of publishing, degree of professionalization, and attitudes toward and engagement with LLMs. Through a grounded theory analysis, we report on core principles that writers express and how these principles can be at odds with their realistic expectations for how institutions engage with their work.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing HCI theory, concepts and modelsccs: Computing methodologies Natural language generation
@@ 1. Introduction
As large language models have gained wide-spread usership, so have investigations into the text used to train them (Chang et al., 2023; Gururangan et al., 2022a; Luccioni and Viviano, 2021; Dodge et al., 2021; Gehman et al., 2020; Bandy and Vincent, 2021; Longpre et al., 2024a). There has been a particularly swift and negative response to the use of creative writing, such as novels, memoirs, and longform journalistic books, as training data, as the authors of this writing express outrage at the use of their work without their consent and without any compensation (Asmelash, 2023). Such frustration has led to dozens of lawsuits against technology companies like OpenAI, Microsoft, and Google, alleging that the use of this writing goes against U.S. copyright law (gen, [n. d.]).
@@
In this work, we focus on understanding the response of creative writers such as novelists, memoirists, poets, and fan fiction writers, all of whom engage in book-length writing that has been used as LLM training data without their consent. While focusing on creative writing may seem niche in the larger world of textual data collection, book-length writing has been found experimentally to be amongst the most highly valuable training data for LLMs (Longpre et al., 2024c). And, to the best of our knowledge, while there have been a series of licensing agreements between technology corporations and news publications (e.g., (O’Brien, 2023)) there is no LLM available that has been trained on consensually collected writing of creative writers. Information scholars have argued that web-scale data collection practices challenge traditional ethical principles governing the use of human subjects data, and that data should be collected and used in accordance with data subjects’ perspectives (Vitak et al., 2016), necessitating research into the perspectives of those creating valuable data. Additionally, as the web becomes increasingly hostile to scrapers (Longpre et al., 2024b) and content is locked down by AI companies (Anna Tong and Coulter, 2024), the need to understand what motivates—or demotivates—people to participate in consensual sharing becomes important to permitting an open web.
With all this playing out in the public and legal sphere, we see a place for researchers to characterize the concerns of creative writers and explore how researchers, technologists, and digital artists may appropriately make use of creative writing to train language models, if at all. In this work, we ask:
@@
To answer these questions, we conducted a grounded theory investigation by interviewing 33 creative writers. We sought to interview a wide variety of authors, in terms of genre, delivery method (i.e., method of sharing or publishing their work), degree of professionalization, demographics, and engagement with AI in their writing practice. We did not presuppose that writers should, or should not, want their writing used as training data. Instead, we sought to understand how they are reasoning about this issue, and the concerns that they raise.
Through a grounded theory analysis, we found that writers reason about the use of their writing as training data according to three interconnected principles. 1) Writers understand that they learn from and influence each other via the creative chain, and question what role LLMs play in this typical process. 2) Writers desire respect for their work as time-consuming labor that requires expertise. 3) Writers note that the human element of writing, particularly embodiment and emotion, means that AI cannot replicate writing processes.
These principles were often at odds with writers’ realistic expectations of the world they live in. 1) Writers experience a lack of control over how their writing is used by technology companies, and how LLMs that may be trained on their writing are used more generally. 2) Writers struggle to make sense of their contribution to LLMs given their interpretation of scale, when any individual’s contribution is so small. 3) Writers have both current and predicted concerns about how this technology will have career impacts and impacts on their industry more broadly.
In the discussion, we return to prior work on archival practices for data collection in machine learning contexts, and note how our findings point towards new principles for data collection in this context. We then address how LLMs challenge writers’ notion of participating in a larger cultural project, in particular how to navigate the decontextualization that LLMs create. Finally, we outline what kind of research might best support writing communities.
@@ 2. Background
In parallel, concerns about the legality of using published books as training data are being raised. The U.S. Copyright office started a study in 2023 regarding copyright issues raised by AI, including several requests for public comment (Office, [n. d.]). Dozens of lawsuits against technology corporations were filed, disputing the use of published books as training data, for the most part due to violations of copyright (gen, [n. d.]). Such investigations and disputes are ongoing.
@@ 3.1. AI Writing Assistance
Computational processes have a long history in creative writing. Early explorations include the surrealistic Dadaism movement and the French Oulipo group, both of which engaged in algorithmic processes for creative writing. In a review of all kinds of writing assistants from 2024, it was found that in the HCI and NLP research communities, papers on writing assistants have been growing dramatically since the mid-2010s (Lee et al., 2024). Recently, much of this work investigates how LLMs, particularly large-scale corporate LLMs like GPT-4, might contribute to particular creative writing genres, like playwriting (Mirowski et al., 2023), storytelling (Ippolito et al., 2022), and cartoon-caption (Kariyawasam et al., 2024), or to particular parts of the creative writing process, like character development (Qin et al., 2024), worldbuilding (Chung and Kreminski, 2024), and revision (Benharrak et al., 2024). Such work demonstrates the capabilities of LLMs to contribute to creative writing, but does little to address why such a system may or may not be adopted by writers. Nor does this work address the documented negative attitude that many creative writers have towards LLMs more generally.
@@ 3.2. Ethical Considerations of Data Usage
Prior to the current interest in generative AI, researchers investigated the ethics of using social media data in HCI and the computational social sciences. Like LLM training data, social media data repurposes people’s data in ways they may not have expected and at a scale that challenges traditional ethical principles governing the ethical use of data involving human subjects. Informed consent, for example, is often infeasible on the scale of social media datasets. Instead, information scholars argue that data should be collected and used in accordance with data subjects’ perspectives on the acceptable use of their data (Hemphill et al., 2022; Fiesler and Proferes, 2018; Vitak et al., 2016).
Research on social media data suggests that data subjects’ level of comfort with their data’s use depends on contextual factors such as the sensitivity of the data, who the researchers are, or what the purpose of the data analysis is. (Hemphill et al., 2022; Fiesler and Proferes, 2018; Gilbert et al., 2021, 2023). For example, Klassen et al. (2022) investigate the ethics of research on Twitter data posted by Black people, finding that several participants expressed that White people doing research on specifically Black Twitter seemed colonizing, whereas the study of Black Twitter was not inherently problematic on its own. These findings suggest that writers may also be concerned about the intentions of the people collecting training data and training LLMs, as opposed to being against the use of their work as training data in all cases.
@@
Just as writers are grappling with the use of their writing as training data, visual artists too are expressing concern, distaste, and often outrage at the use of visual art as training data for generative AI, specifically text-to-image models. Artists have launched a number of their own lawsuits against technology companies (gen, [n. d.]), often analogous to those from the writing community. Research has articulated different kinds of harm AI art is creating for artists, including direct economic harms as well as chilling effects (that is, generative AI is discouraging artists from sharing their work online, making it harder for artists to find their audience) (Jiang et al., 2023). Lovato et al. (2024), in a survey of over 400 visual artists, found that a majority of artists believe training data for generative models should be disclosed, and express concerns about the impact of AI on their industry (Lovato et al., 2024).
@@ 3.3. Addressing Concerns about Data Usage
Technical research attempts to address artists’ concerns about training data. To address concerns that generative AI may plagiarize pretraining data contributors, researchers have proposed approaches to minimize model memorization (Ippolito et al., 2023; Carlini et al., 2023; Kandpal et al., 2022; Lee et al., 2022a; Vyas et al., 2023; Chu et al., 2024) and style mimicry (Shan et al., 2023; Liang et al., 2023; Van Le et al., 2023). Another line of work attempts to protect the IP of writers or programmers by watermarking text (Lau et al., 2024; Qiang et al., 2023; Yoo et al., 2023) and attributing generated content to specific training data (Deng et al., 2024; Hammoudeh and Lowd, 2024). Still another proposed approach is to train language models on permissively licensed text, and then query copyrighted materials in a datastore, allowing content creators to opt-out and remove their material from the datastore without needing to retrain the model (Min et al., 2023). To enable individuals to opt out of training data, researchers propose methods for unlearning (Chen and Yang, 2023; Jang et al., 2023) and modular language model architectures (Gururangan et al., 2022b; Li et al., 2022). Such work presents potential technical solutions, but does not answer questions about whose work should and should not be protected and under what contexts.
@@ 4.1. Procedure and Analysis
Then, axial and selective coding was done on all codes that came out of the open coding performed on each index code. This allowed us to see emergent and general themes that were occurring across the specific concerns that writers were bringing up. With these new, general codes, we performed a final round of coding in which all excerpts were re-coded with the general codes to both test their performance on the data, and to note any data that did not fit into this framework. These new, general codes are what are reported as our results.
@@ 4.2. Recruitment
Although we did not recruit for this specifically, we found that our participants had a wide range of views on language models more generally. Some considered the technology to have no positive applications (and thus could not think of any reason to contribute to them), others considered the technology to be neutral, and still others thought the technology had pro-social applications, even if they were against the current way in which they were created or used.
@@ 5. Findings
First, we report on the principles that are the foundation for how writers think about writing in general, and that they use to reason about the use of their writing in this specific context. Second, we report on their realistic expectations about how the world works. Here, we see writers’ principles can sometimes, but not always, clash against what they think is possible in the world they live in.
We note that there are a variety of ideas that thread throughout multiple codes. For instance, a major discussion point was if and how compensation may work. This idea comes up under the code respect (where compensation is often seen as a mechanism of showing respect) as well as under the code interpretation of scale (where writers reason about how much compensation could reasonably be expected) and lack of control (where writers note that it may be impossible to control how their writing is used, negating the possibly for fair compensation). Our codes outline how writers are reasoning about these issues, rather than presenting a list of concerns.
@@ 5.1.1. The Creative Chain
Participants often compared making their writing accessible to train a language model to their principles and expectations around making their writing accessible in other contexts. In general, participants recognize that reading, imitating, referencing and transforming other writing is essential both to learning to write and writing itself. In turn, many participants described a responsibility to support education and the production of art by making their writing accessible and taking a permissive stance towards work that references theirs. As P12 put it, “I think that being very rigid around intellectual property, and the defense of ideas… turns into an awful scenario that I don’t really believe in, because I believe that all art inspires other art, and I love that basic premise.” P20 described this as “being a link in the creative chain”: “We’re all building off of each other constantly, you know. My work is just the product of everything I’ve absorbed through my life… nothing makes me happier than seeing people make other things with my work.
@@
Making information free and accessible for others to learn from is an essential principle for several participants. Because of this, some writers felt ambivalent or positive about their writing being used to train language models, despite other concerns they might have (P5, P27, P24, P25, P17, P30).
@@
Expectations of control in the creative chain: What is the project I am a part of? In contrast to other creative projects a writer might contribute their work to, some participants expressed that the overall goal of language models was unclear to them, and it was therefore difficult to assess what it meant to contribute their writing. P3 noted that “it feels similar to what I imagine an artist who’s being curated might feel: … they would want to know about the content or subject, or the nature of the thing that they’re being curated into.” Understanding the goal of the project was important for P14, who said, “If I knew the end result or kind of, Here’s why we’re interested in your stuff, I would be much more interested in opting in.
@@
Disagreements about the artistic potential of language models. Participants disagreed about whether contributing to training datasets could yield benefits such as supporting the production of worthwhile art, helping others to express themselves, or supporting education. Several participants had trouble imagining the artistic potential of language models because they did not like the style of generated text (P21, P25), finding it “generic” (P13, P6, P24), or had concerns about the political biases of these models (P21, P18). Many others expressed doubts about the creative potential of LLMs, both because they view LLMs as “regurgitation” (P18) of other work rather than creating new ideas, and because of the fundamental role of humanity in creativity. (See subsubsection 5.1.3.)
@@
On the other hand, several writers reported regularly using LLMs in their creative process (P1, P8, P15, P28), or cited literary work that made use of language models that they liked (P22, P4, P12, P15). Many others saw the value in using LLMs “as a tool” (P2) for creative expression, imagining that LLMs might help writers generate initial ideas (P27, P18), provide writers with feedback or analytic summaries during the writing process (P3, P27, P5, P16), or help writers move past writer’s block (P15, P18, P13, P3, P29).
@@
Representation and nourishing the future of technology. In addition to supporting the production of art, writers identified other positive, meaningful aspects of contributing their writing to training data. Some writers expressed excitement about the idea of “contribut[ing] to the future” (P28) or “nourish[ing]” (P15) language models with literary writing (P1, P26). As P11 explained, “I’ve had an influence on the way this machine thinks, which is incredible. I mean, it’s probably that’s my place in history more than any encyclopedia entry, you know.
@@
Some writers with racial identities that are under-represented both in language model outputs and in the literary world more broadly were enthusiastic about contributing their data (P27, P23), as “there needs to be more diversity in some of these models” (P27).
@@
Decontextualization: being included could misrepresent or reify a writer’s style, work, and identity. Some participants expressed concerns about how contributing to language models had the potential to decontextualize their work. For P24, decontextualization is a threat to their artistic evolution and identity: “I don’t want my past books to be training… because I want to always evolve and become a better writer, and I feel like that would leave me stuck in that same voice or ability.” P32 also expressed concerns about decontextualization, as they thought through what would come up if someone prompted a model trained on writing they wrote throughout their gender transitions. They were comfortable with someone reading their old work online, where it would be accompanied by contextual information like when it was written (”I trust readers, I think readers are really smart”); in an LLM, they feared it could instead be weaponized against their community.
@@
For P19, the way language models racially decontextualize their training data is a big concern:
@@
Expectations of respect and humanity in the creative chain. Participants made analogies between contributing their writing to language model training datasets and various ways their work does or could relate to other creative projects: it is like someone turning your book into a movie or play (P3, P20, P12), it is like having your style mimicked by your student or competition (P31, P7, P6), it is like having your words appear uncredited in someone else’s work (P31, P13, P18, P12, P29, P17), it is like someone writing fanfiction about your novel (P1, P18, P20), it is like being curated into someone’s art show (P3). Participants described how these different scenarios present different risks to their careers, reputations, and sense of creative control. As such, in each of these scenarios, participants described how they expect to receive respect through some combination of credit, compensation, creative control, or at least the sense that a human being who admires their writing engaged with it, expressing a shared value of creative writing.

### v2

@@ Abstract.
The use of creative writing as training data for large language models (LLMs) is highly contentious and many writers have expressed outrage at the use of their work without consent or compensation. In this paper, we seek to understand how creative writers reason about the real or hypothetical use of their writing as training data. We interviewed 33 writers with variation across genre, method of publishing, degree of professionalization, and attitudes toward and engagement with LLMs. We report on core principles that writers express (support of the creative chain, respect for writers and writing, and the human element of creativity) and how these principles can be at odds with their realistic expectations of the world (a lack of control, industry-scale impacts, and interpretation of scale). Collectively these findings demonstrate that writers have a nuanced understanding of LLMs and are more concerned with power imbalances than the technology itself.
@@
journalyear: 2025copyright: acmlicensedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713287isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing HCI theory, concepts and modelsccs: Computing methodologies Natural language generation
@@ 1. Introduction
As large language models have gained wide-spread usership, so have investigations into the text used to train them (Chang et al., 2023; Gururangan et al., 2022a; Luccioni and Viviano, 2021; Dodge et al., 2021; Gehman et al., 2020; Bandy and Vincent, 2021; Longpre et al., 2024a). There has been a particularly swift and negative response to the use of creative writing, such as novels, memoirs, and longform journalistic books, as training data, as the authors of this writing express outrage at the use of their work without consent or compensation (Asmelash, 2023). Such frustration has led to dozens of lawsuits against technology companies like OpenAI, Microsoft, and Google, alleging that this use of writing goes against U.S. copyright law (gen, [n. d.]).
@@
In this work, we focus on understanding the response of creative writers such as novelists, memoirists, poets, and fan fiction writers, all of whom engage in book-length writing that has been used as LLM training data without their consent. While focusing on creative writing may seem niche in the larger world of textual data collection, book-length writing has been found experimentally to be among the most highly valuable training data for LLMs (Longpre et al., 2024c). And, to the best of our knowledge, while there have been a series of licensing agreements between technology corporations and news publications (e.g., (O’Brien, 2023)) there is no LLM available that has been trained on the consensually collected writing of creative writers. Information scholars have argued that web-scale data collection practices challenge traditional ethical principles governing the use of human subjects data, and that data should be collected and used in accordance with data subjects’ perspectives (Vitak et al., 2016), necessitating research into the perspectives of those creating valuable data. Additionally, as the web becomes increasingly hostile to scrapers (Longpre et al., 2024b) and content is locked down by AI companies (Anna Tong and Coulter, 2024), the need to understand what motivates—or demotivates—people to participate in consensual sharing becomes important to permitting an open web.
With all this playing out in the public and legal sphere, we see a place for researchers to characterize the concerns of creative writers. We do not investigate these issues from a legal perspective, but rather an ethical and human-centered one, centering the voices of creative writers. In this work, we ask:
@@
To answer these questions, we conducted a grounded theory investigation by interviewing 33 creative writers. We sought to interview a wide variety of authors, in terms of genre, delivery method (i.e., method of sharing or publishing their work), degree of professionalization, demographics, and engagement with AI in their writing practice. We did not presuppose that writers should, or should not, want their writing used as training data. Instead, we sought to understand how they are reasoning about this issue and the concerns that they raise.
We found that writers reason about the use of their writing as training data according to three interconnected principles. 1) Writers understand that they learn from and influence each other via the creative chain, and question what role LLMs play in this process. 2) Writers desire respect for their work as time-consuming labor that requires expertise. 3) Writers note that the human element of writing, particularly embodiment and emotion, means that AI cannot replicate writing processes.
These principles were often at odds with writers’ realistic expectations of the world they live in. 1) Writers experience a lack of control over how their writing is used by technology companies, and how LLMs that may be trained on their writing are used more generally. 2) Writers struggle to make sense of their contribution to LLMs given their interpretation of scale, when any individual’s contribution is so small. 3) Writers have both current and predicted concerns about how this technology will have industry impacts, some of which supersede personal impacts.
In the discussion, we consider how writing changes when it becomes data, and propose a colonial lens to understand paths forward. We also discuss the place of LLMs in the creative chain, and how they may (or may not) be like libraries. Finally, we outline what kind of research might best support writing communities.
@@ 2. Background
As writers express their discontent, concerns about the legality of using published books as training data are being raised. The U.S. Copyright office started a study in 2023 regarding copyright issues raised by AI, including several requests for public comment (Office, [n. d.]). Dozens of lawsuits against technology corporations were filed, disputing the use of published books as training data, for the most part due to violations of copyright (gen, [n. d.]). Such investigations and disputes are ongoing. Legal questions will be resolved by the courts, and likely differ across jurisdictions and nations. In our work, we seek to uncover not strong legal arguments, but rather the humanistic concerns and reasoning of writers.
@@ 3.1. AI Writing Assistance
Computational processes have a long history in creative writing. Early explorations include the surrealistic Dadaism movement and the French Oulipo group, both of which engaged in algorithmic processes for writing. A 2024 review of all kinds of writing assistants found that, in the HCI and NLP research communities, papers on writing assistants have been growing dramatically since the mid-2010s (Lee et al., 2024). Recently, much of this work investigates how LLMs, particularly large-scale corporate LLMs like GPT-4, might contribute to particular creative writing genres, like playwriting (Mirowski et al., 2023), storytelling (Ippolito et al., 2022), and cartoon-captioning (Kariyawasam et al., 2024), or to particular parts of the creative writing process, like character development (Qin et al., 2024), worldbuilding (Chung and Kreminski, 2024), and revision (Benharrak et al., 2024). Such work demonstrates the capabilities of LLMs to contribute to creative writing, but does little to address why such a system may or may not be adopted by writers. Nor does it address the documented negative attitude that many creative writers have towards LLMs more generally.
@@ 3.2. Ethical Considerations of Data Usage
Prior to the current interest in generative AI, researchers investigated the ethics of using social media data in HCI and the computational social sciences. Like LLM training data, social media data repurposes people’s data in ways they may not have expected and at a scale that challenges traditional principles governing the ethical use of data involving human subjects. Informed consent, for example, is often infeasible on the scale of social media datasets. Instead, information scholars argue that data should be collected and used in accordance with data subjects’ perspectives on the acceptable use of their data (Hemphill et al., 2022; Fiesler and Proferes, 2018; Vitak et al., 2016).
Research on social media data suggests that data subjects’ level of comfort with their data’s use depends on contextual factors such as the sensitivity of the data, who the researchers are, or what the purpose of the data analysis is (Hemphill et al., 2022; Fiesler and Proferes, 2018; Gilbert et al., 2021, 2023). For example, Klassen et al. (2022) investigate the ethics of research on Twitter data posted by Black people, finding that several participants expressed that White people doing research on specifically Black Twitter seemed colonizing, whereas the study of Black Twitter was not inherently problematic on its own. These findings suggest that writers may also be concerned about the intentions of the people collecting training data and training LLMs, as opposed to being against the use of their work as training data in all cases.
@@
Just as writers are grappling with the use of their writing as training data, visual artists too are expressing concern, distaste, and often outrage at the use of visual art as training data for generative AI, specifically text-to-image models (Jiang et al., 2023). Artists have launched a number of their own lawsuits against technology companies (gen, [n. d.]), often analogous to those from the writing community. Research has articulated different kinds of harm AI art is creating for artists, including direct economic harms as well as chilling effects (that is, generative AI is discouraging artists from sharing their work online, making it harder for artists to find their audience) (Jiang et al., 2023). Lovato et al. (2024), in a survey of over 400 visual artists, found that a majority of artists believe training data for generative models should be disclosed and express concerns about the impact of AI on their industry (Lovato et al., 2024). We see this work as a primer for our own, as creative writers may or may not reason about generative AI in the same way as visual artists, and individual and industry impacts may differ.
@@ 3.3. Addressing Concerns about Data Usage
Technical research attempts to address artists’ concerns about training data. To address concerns that generative AI may plagiarize pretraining data contributors, researchers have proposed approaches to minimize model memorization (Ippolito et al., 2023; Carlini et al., 2023; Kandpal et al., 2022; Lee et al., 2022a; Vyas et al., 2023; Chu et al., 2024) and style mimicry (Shan et al., 2023; Liang et al., 2023; Van Le et al., 2023). Another line of work attempts to protect the IP of writers or programmers by watermarking text (Lau et al., 2024; Qiang et al., 2023; Yoo et al., 2023) and attributing generated content to specific training data (Deng et al., 2024; Hammoudeh and Lowd, 2024). To enable individuals to opt out of training data, researchers propose methods for unlearning (Chen and Yang, 2023; Jang et al., 2023) and modular language model architectures (Gururangan et al., 2022b; Li et al., 2022; Min et al., 2023). Such work presents potential technical solutions, but does not answer questions about whose work should be protected and under what contexts.
@@ 4.1. Procedure and Analysis
Then, axial and selective coding was done on all codes that came out of the open coding performed on each index code. This allowed us to see emergent and general themes that were occurring across the concerns that writers were bringing up. With these new, general codes, we performed a final round of coding in which all excerpts were re-coded with the general codes to both test their performance on the data, and to note any data that did not fit into this framework. These codes are what are reported as our results.
@@ 4.2. Recruitment
Although we did not recruit for this specifically, we found that our participants had a wide range of views on language models more generally. Some considered the technology to have no positive applications, others considered it to be neutral, and still others thought it had pro-social applications, even if they were against the current way in which it was created or used.
@@ 5. Findings
All our codes can be found in Table 1. First, we report on the principles that are the foundation for how writers think about writing in general, and that they use to reason about the use of their writing in this new context. Second, we report on their realistic expectations about how the world works. Here, we see writers’ principles can sometimes, but not always, clash against what they think is possible in the world they live in.
We note that there are a variety of ideas that thread throughout multiple codes. For instance, a major discussion point was if and how compensation may work. This idea comes up under the code respect (where compensation is often seen as a mechanism of showing respect) as well as under the code interpretation of scale (where writers reason about how much compensation could reasonably be expected) and lack of control (where writers note that it may be impossible to control how their writing is used, negating the possibility for fair compensation). Our codes outline how writers are reasoning about these issues, rather than presenting a list of concerns.
Table 1. Overview of the codes that emerged from our grounded theory analysis.
Principles
The Creative Chain It’s just being a link in the creative chain. I made something. Somebody else saw that, and they got a spark of an idea to do something else, and they want to build off of it. We’re all building off of each other constantly.” (P20)
Respect I think it just comes again to respecting the works and respecting the people. Because I think that’s probably my main issue with a lot of AI stuff is just the complete lack of respect for the works that they have been using to profit off.” (P18)
The Human Element Honestly, I think a lot of the reason that people like reading other people’s stuff is because there’s the human element there. You’re looking at one guy’s perspective of an insane situation … It’s the project of one person’s mind.” (P13)
Realistic Expectations
Lack of Control One of the things that we’re all really feeling right now is a lack of any options in this. When I found out that my work was in Books3 I emailed my agent right away. And I was like, I know there’s like nothing you can do, I just want you to know that this is a thing, and that I don’t like it.” (P20)
Industry Impacts We could easily be putting artists out of work. And what does that mean for them as artists and humans, [who need] to put food on their table, but also like, what does that mean for culture and society?” (P9)
Interpretation of Scale Cause I’m trying to think, okay, upsides and downsides. It’s hard to pinpoint one, because again, I’m a little drop of water in a huge ocean. I don’t have a distinct style, so I don’t… I don’t really know.” (P24)
@@ 5.1.1. The Creative Chain
Participants often compared making their writing accessible to train a language model to their principles and expectations around making their writing accessible in other contexts. In general, participants recognize that reading, imitating, referencing and transforming other writing is essential both to learning to write and writing itself. In turn, many participants described a responsibility to support education and the production of art by making their writing accessible and taking a permissive stance towards work that references theirs. As P12 (professional novelist) put it, “I think that being very rigid around intellectual property, and the defense of ideas… turns into an awful scenario that I don’t really believe in, because I believe that all art inspires other art, and I love that basic premise.” P20 (professional novelist) described this as “being a link in the creative chain”: “We’re all building off of each other constantly, you know. My work is just the product of everything I’ve absorbed through my life… nothing makes me happier than seeing people make other things with my work.
@@
Making information free and accessible for others to learn from was an essential principle for several participants. Because of this, some writers felt ambivalent or positive about their writing being used to train language models, despite other concerns they might have (P5, P17, P24, P25,P27, P30).
@@
Expectations of control in the creative chain: What is the project I am a part of? In contrast to other creative projects a writer might contribute their work to, some participants expressed that the overall goal of language models was unclear to them, and it was therefore difficult to assess what it meant to contribute their writing. P3 (poet and essayist) noted that “it feels similar to what I imagine an artist who’s being curated might feel: … they would want to know about the content or subject, or the nature of the thing that they’re being curated into.” Understanding the goal of the project was important for P14 (poet), who said, “If I knew the end result or kind of, Here’s why we’re interested in your stuff, I would be much more interested in opting in.
@@
Disagreements about the artistic potential of language models. Participants disagreed about whether contributing to training datasets could yield benefits such as supporting the production of worthwhile art, helping others to express themselves, or supporting education. Several participants had trouble imagining the artistic potential of language models because they did not like the style of generated text (P21, P25), found it “generic” (P6, P13, P24), or had concerns about the political biases of these models (P18, P21). Many others expressed doubts about the creative potential of LLMs, both because they view LLMs as “regurgitation” (P18, fan fiction writer) of other work rather than creating new ideas, and because of the fundamental role of humanity in creativity. (See the human element 5.1.3.)
@@
On the other hand, several writers reported regularly using LLMs in their creative process (P1, P8, P15, P28), or cited literary work that used language models that they liked (P4, P12, P15, P22). Many others saw the value in using LLMs “as a tool” (P2) for creative expression, imagining that LLMs might help writers generate initial ideas (P18, P27), provide writers with feedback or analytic summaries during the writing process (P3, P5, P16, P27), or help writers move past writer’s block (P3, P13, P15, P18, P29).
@@
Representation and nourishing the future of technology. In addition to supporting the production of art, writers identified other positive, meaningful aspects of contributing their writing to training data. Some writers expressed excitement about the idea of “contribut[ing] to the future” (P28, poet) or “nourish[ing]” (P15, novelist and journalist) language models with literary writing (also P1, P26). As P11, a professional novelist who had never used an LLM, explained, “I’ve had an influence on the way this machine thinks, which is incredible. I mean, it’s probably that’s my place in history more than any encyclopedia entry, you know.
@@
Some writers with racial identities that are under-represented both in language model outputs and in the literary world more broadly were enthusiastic about contributing their data (P23, P27), as “there needs to be more diversity in some of these models” (P27).
@@
Decontextualization: being included could misrepresent or reify a writer’s style, work, and identity. Some participants expressed concerns about how contributing to language models had the potential to decontextualize their work. For P24 (novelist), decontextualization is a threat to their artistic evolution and identity: “I don’t want my past books to be training… because I want to always evolve and become a better writer, and I feel like that would leave me stuck in that same voice or ability.” P32 (professional essayist and novelist) also expressed concerns about decontextualization, as they thought through what would come up if someone prompted a model trained on writing they wrote throughout their gender transitions. They were comfortable with someone reading their old work online, where it would be accompanied by contextual information like when it was written (”I trust readers, I think readers are really smart”); in an LLM, they feared it could instead be weaponized against their community.
@@
For P19, the way language models racially decontextualize their training data was a big concern:
@@
Expectations of respect and humanity in the creative chain. Participants made analogies between contributing their writing to language model training datasets and various ways their work does or could relate to other creative projects: it is like someone turning your book into a movie or play (P3, P12, P20), it is like having your style mimicked by your student or competition (P6, P7, P31), it is like having your words appear uncredited in someone else’s work (P12, P13, P17, P18, P29, P31), it is like someone writing fanfiction about your novel (P1, P18, P20), it is like being curated into someone’s art show (P3). Participants described how these different scenarios present different risks to their careers, reputations, and sense of creative control. As such, in each of these scenarios, participants described how they expect to receive respect through some combination of credit, compensation, creative control, or at least the sense that a human being who admires their writing engaged with it, expressing a shared value of creative writing.